# On the randomized Euler scheme for SDEs with integral-form drift

### v1

@@ 7. Appendix
    Let $H\in\mathcal{A}(L,K,p),$ then
    $\|a(x)\|\leq\max\{\|H(\xi,0)\|_{p},\|L\|_{L^{P}(\mathcal{T},\mathcal{A},\mu)}\}(1+\|x\|),$
@@
For all $n\in\mathbb{N}$ and $(M_{k})_{k=1}^{n}\subset\mathbb{N},$ it holds that
@@
since $\tilde{X}_{n,M}^{RE}$ is a linear interpolation of $X_{n,M}^{RE}.$ For all $n\in\mathbb{N},(M_{k})_{k=1}^{n}\subset\mathbb{N}$ and $k=0,\ldots,n-1,$ random variables $b(X_{n,M}^{RE}(t_{k}))$ and $\Delta W_{k}$ are independent. Henceforth,
$\begin{split}\|X_{n,M}^{RE}(t_{k+1})\|_{p}&\leq\|X_{n,M}^{RE}(t_{k})\|_{p}+\frac{h}{M_{k+1}}\sum\limits_{j=1}^{M_{k+1}}\|H(\xi_{j}^{k+1},X_{n,M}^{RE}(t_{k}))\|_{p}+\|b(X_{n,M}^{RE}(t_{k}))\Delta W_{k}\|_{p}\\ &=\|X_{n,M}^{RE}(t_{k})\|_{p}+h\|H(\xi,X_{n,M}^{RE}(t_{k}))\|_{p}+\|b(X_{n,M}^{RE}(t_{k}))\Delta W_{k}\|_{p}\\ &\leq\|X_{n,M}^{RE}(t_{k})\|_{p}+T\|H(\xi,X_{n,M}^{RE}(t_{k}))\|_{p}+\|b(X_{n,M}^{RE}(t_{k}))\|_{p}\|\Delta W_{k}\|_{p}.\end{split}$
Finally, the thesis follows immediately from the aforementioned observation, mathematical induction, and at-most linear growth of $H$ and $b$ from Fact 1. ∎
@@
There exists constant $\tilde{C}_{p}\in[0,+\infty)$ depending only on $\eta,T$ and parameters of the class $\mathcal{F}(L,K,p),$ such that for all $(a,b)\in\mathcal{F}(L,K,p),M,n\in\mathbb{N}$ and all $k=0,1,...,n-1$ the following inequality holds
@@
Since function $[0,T]\ni t\mapsto\sup\limits_{0\leq u\leq t}\mathbb{E}\|\tilde{X}_{n,M}^{RE}(u)\|^{p}$ is Borel-measurable (as a non-decreasing function) and bounded (by fact 2), applying the Gronwall’s lemma yields
@@
Acknowledgements We gratefully acknowledge the support of the Leibniz Center for Informatics, where several discussions about this research were held during the Dagstuhl Seminar "Algorithms and Complexity for Continuous Problems" (Seminar ID 23351).

### v2

@@ 7. Appendix
    Let $H\in\mathcal{A}(L,K,p)$ and let $a(x)=\mathbb{E}(H(\xi,x)),$ then
    $\|a(x)\|\leq\max\{\|H(\xi,0)\|_{p},\|L\|_{L^{P}(\mathcal{T},\mathcal{M},\mu)}\}(1+\|x\|),$
@@
For all $n\in\mathbb{N}$ and $M\in\mathbb{N}$,it holds that
@@
since $\tilde{X}_{n,M}^{RE}$ is a linear interpolation of $X_{n,M}^{RE}.$ For all $n,M\in\mathbb{N}$ and $k=0,\ldots,n-1,$ random variables $b(X_{n,M}^{RE}(t_{k}))$ and $\Delta W_{k}$ are independent. Henceforth,
$\begin{split}\|X_{n,M}^{RE}(t_{k+1})\|_{p}&\leq\|X_{n,M}^{RE}(t_{k})\|_{p}+\frac{h}{M}\sum\limits_{j=1}^{M}\|H(\xi_{j}^{k+1},X_{n,M}^{RE}(t_{k}))\|_{p}+\|b(X_{n,M}^{RE}(t_{k}))\Delta W_{k}\|_{p}\\ &=\|X_{n,M}^{RE}(t_{k})\|_{p}+h\|H(\xi,X_{n,M}^{RE}(t_{k}))\|_{p}+\|b(X_{n,M}^{RE}(t_{k}))\Delta W_{k}\|_{p}\\ &\leq\|X_{n,M}^{RE}(t_{k})\|_{p}+T\|H(\xi,X_{n,M}^{RE}(t_{k}))\|_{p}+\|b(X_{n,M}^{RE}(t_{k}))\|_{p}\|\Delta W_{k}\|_{p}.\end{split}$
Finally, the thesis follows immediately from the aforementioned observation, mathematical induction, and at most linear growth of $H$ and $b$ from Fact 1. ∎
@@
There exists constant $\tilde{C}_{p}\in[0,+\infty)$ depending only on $\eta,T$ and parameters of the class $\mathcal{F}(L,K,p),$ such that for all $(H,b)\in\mathcal{F}(L,K,p),M,n\in\mathbb{N}$ and all k=0,1,,n1k=0,1,...,n-1 the following inequality holds
@@
Since function $0,T]\ni t\mapsto\sup\limits_{0\leq u\leq t}\mathbb{E}\|\tilde{X}_{n,M}^{RE}(u)\|^{p}$ is Borel-measurable (as a non-decreasing function) and bounded (by fact 2), applying the Gronwall’s lemma yields
@@
Acknowledgements We gratefully acknowledge the support of the Leibniz Center for Informatics, where several discussions about this research were held during the Dagstuhl Seminar "Algorithms and Complexity for Continuous Problems" (Seminar ID 23351). We want to thank the unknown reviewers for their valuable comments and suggestions on improving the overall quality of this work.
Statements and Declarations
The authors declare no competing interests. No funding was received to assist with the preparation of this manuscript. The datasets generated during and/or analysed during the current study are available from the authors on reasonable request.